OpenStudy (anonymous):

6/12 - 1/3 in simplest form i'm dumb :/

OpenStudy (anonymous):

1/2-1/3

OpenStudy (anonymous):

6/12 - 1/3 = 6/12 - 4/12 = 2/12 = 1/6 <-- Simplest form.

OpenStudy (anonymous):

it would be 1/2

OpenStudy (anonymous):

@genius12 THX!!!!!!!!!!